# Probing Langmuir monolayer self-assembly in condensed and collapsed phases: grazing incidence X-ray diffraction and X-ray standing waves studies

### v1

@@ 4 Conclusions
We investigated the effect of aqueous Ce3+ ions on the structural reorganization of arachidic acid monolayers at different temperatures and surface pressures. The morphology of the monolayer was observed by BAM. The structures were deduced from synchrotron measurements: the lateral ordering was studied with the GID and the vertical distribution of Ce ions with the XSW. By varying the surface pressure and temperature, we observed the structural reorganization of the system. In these observations we reproduced a number of known Ce-induced effects on fatty acid self-assembly. First, the binding of the metal to the carboxylic acid groups leads to the formation of a rigid monolayer with higher crystallinity and reduced molecular area, and such solid-phase monolayers are formed during compression by bypassing the liquid phase, as evidenced by the - isotherm measurements.
Second, by interacting with the carboxylic acid groups, the Ce ions may form a laterally ordered layer underneath the organic monolayer. In the GID measurements we observed an inorganic superstructure of Ce ions which is commensurate with the fatty acid monolayer having a unit cell. The localization of this Ce layer underneath the fatty acid monolayer is confirmed by the XSW measurements. The formation of such an inorganic Ce superstructure can be reproducibly observed at C.
The third Ce-induced effect was observed in the solid-phase monolayers formed at a lower temperature: C. Initially, the system behaves like a solid phase at C. However, with time, another superstructure appears: an organic superstructure. This superstructure does not consist of Ce ions. Instead, it is formed by the molecules of a fatty acid itself. At lower temperatures, the energy stored in the rotational motion of the molecules is reduced, allowing the mutual orientation of adjacent aliphatic chains, which in turn allows the molecular area to be further reduced, forming an even denser monolayer. When the angle between the aliphatic chains of adjacent molecules is about , such a configuration is called PHB. This is an organic superstructure with cells. Consequently, this breaks the symmetry in the structure factor and an additional peak appears in the GID pattern, which we actually observed. Furthermore, the crystal structure we deduced from the GID pattern corresponds to the PHB phase mode according to the literature.
Finally, we observed a new collapse mode of AA monolayers. Surprisingly, in this collapse mode the monolayer maintains its translational order despite being compressed beyond the collapse point. Instead, the collapse is manifested by the distortion of the long-range orientational order. This is hinted at in BAM images, where a mosaic-like morphology of the film can be seen in the new collapse mode. This is confirmed by GID. Interpretation of the GID data leads to the following conclusions. The monolayer in the new collapse mode consists of 2D crystalline domains that are inclined from the horizontal position. The crystalline structure of a domain corresponds to the PHB packing mode. The domains are inclined according to a normal distribution. We deduced this by fitting forward simulations based on the DWBA to the GID data. Unlike the other collapse modes, the GID pattern did not decay with time. The only change in GID we observed can be attributed to a change in the distribution of domain inclination, the translational order, i.e. crystalline structure of the domains themselves, remained unchanged. By comparing AA monolayers in solid phase and in collapse, we can assume that the packing mode of a solid phase can influence which collapse scenario is realized. Namely, PHB seems to lead to collapse mode 2. Speculatively, the PHB, being a tighter packing than the free-rotating state of the AA monolayer, may explain that the translational order is unchanged in collapse. This is a subject for further study. With this article, we hope to expand the knowledge of the structural reorganization of fatty acids, which may be important for the functionalization of nanomaterials.
@@
Formation of phospholipid monolayer on liquid surface Nitrate salt Ce(NO3)6H2O (99.99%) was purchased from LANHIT Ltd and used as received. Arachidic acid was purchased from Sigma. To form a lipid monolayer, a solution of arachidic acid in chloroform (concentration of 0.59 mg/ml) was spread on the surface of Ce(NO3)6H2O aqueous solution at a solution pH of . The solvent was allowed to evaporate for 15 minutes. The layer was then compressed in a Langmuir trough to the desired surface pressure. All solutions were prepared using ultrapure water Milli-Q Advantage A10 Water Purification System, (Millipore).
Surface preasure-area measurements The Langmuir films were studied on a Minitrough Extended Langmuir trough (KSV) with a maximum interfacial area of 558 cm2. The compression rate was 3.75 cm2/min. External vibrations of the Langmuir trough were damped by an active vibration damper (Accurion).
Brewster Angle Microscopy BAM images were obtained using a Brewster angle microscope (model BAM-300, KSV NIMA) coupled to a temperature-controlled Langmuir trough. A polarized He-Ne laser (wavelength 632.8 nm) was used as the light source, with the beam directed at the air/water interface at an angle of (the Brewster angle for water). The reflected beam was collected with a 10 objective; images were taken with a USB camera and corrected with a KSV NIMA software package. Measurements were performed at 21∘C and 23∘C.
Synchrotron measurements GID and XSW studies were performed at the bending magnet beamline LANGMUIR (Kurchatov Center for Synchrotron Radiation, Moscow, Russia), which is equipped with a customized Langmuir trough. The closed Langmuir trough was flushed with water vapor saturated helium to reduce the scattering of the incident X-rays and to reduce the evaporation of the liquid subphase, thus keeping the liquid surface position fixed. The photon energy of the incident beam for all GID measurements was 13 keV. GID measurements were performed at ( is the critical angle of total external reflection for water). The diffracted intensity in the direction was recorded using a linear position sensitive detector (Mythen) with a vertical stripe size of m. For XSW measurements, the fluorescence signal was recorded by an energy dispersive Vortex EX detector mounted perpendicular to the liquid surface. The characteristic fluorescence spectra were recorded for each angle of incidence in the angular range corresponding to the total external reflection region.
@@ Acknowledgments
The authors are deeply indebted to Dr. Vladimir Kaganer and Dr. Boris Ostrovskii for fruitful discussions, insightful advice, and helpfulness.

### v2

@@ 4 Conclusions
Ce-induced effects on structural reorganization during self-assembly in fatty acid monolayer were investigated depending on temperature. We examined the phase behavior of AA monolayer formed on the aqueous solution of Ce(NO3)6H2O salt. GID and XSW measurements were applied to follow the changes in monolayer packing in condensed and collapsed phases at two different temperatures and . The Ce3+ ions were found to form a laterally ordered layer underneath the solid-phase monolayer in both cases. GID measurements revealed the formation of an inorganic superstructure of Ce3+ ions which is commensurate with the fatty acid monolayer having a superstructure. The localization of the Ce layer underneath the fatty acid monolayer was confirmed by XSW measurements. The structural organization of the fatty acid molecules in the presence of Ce3+ ions at was found to be very typical. In fact, they form a lattice in the solid phase with an NN tilt. The only difference is that the cell is oblique and the packing is tighter as compared to the monolayer on pure water. Standard collapse behavior is also observed for the fatty acid monolayer at . In terms of GID, this is manifested by the broadening of the peaks as the structure degrades. The most pronounced Ce-induced effects on AA self-assembly were observed at . In the condensed phase ( mN/m) we detected the gradual transition from the NN tilted structure to the so-called PHB packing mode – an organic superstructure. At lower temperatures, the energy stored in the rotational motion of the molecules is reduced, allowing the mutual orientation of adjacent aliphatic chains, which in turn allows the molecular area to be further reduced, forming an even denser monolayer.
Surprisingly, in the collapse region at , the monolayer maintains its translational order despite being overcompressed. Quantitative analysis of 2D maps of diffraction scattering provided the fine details of molecular rearrangement in the collapsed AA film. Namely, the monolayer in collapse mode at consists of 2D crystalline flat domains, which are inclined from the horizontal position. The crystalline structure of a domain corresponds to the PHB packing mode. The domains are inclined according to a normal distribution. We deduced this by fitting forward simulations based on the DWBA to the GID data. Importantly, the GID pattern did not decay with time. The observed changes in GID can be attributed to a change in the distribution of domain inclination, while the translational order, i.e. the crystalline structure of the domains themselves, remained unchanged. Note that collapsed structures are exceptionally challenging objects for GID studies. Nevertheless, even in such complicated cases the careful quantitative analysis of the full 2D GID pattern can provide information about fine details of mechanisms governing the collapse processes.
@@
Formation of fatty acid monolayer on liquid surface. Ce(NO3)6H2O (99.99%) was purchased from LANHIT Ltd and used as received. Arachidic acid was purchased from Sigma. To form a monolayer, a solution of arachidic acid in chloroform (concentration of 0.59 mg/ml) was spread on the surface of 20 mM Ce(NO3)6H2O aqueous solution of pH 5.5. Chloroform was allowed to evaporate for 15 minutes, after that the layer was compressed to a desired surface pressure. Cerium nitrate solutions were prepared using ultrapure water (Millipore Milli-Q Advantage A10).
Surface preasure-area measurements. The Langmuir films were studied on a Minitrough Extended Langmuir trough (KSV) with a maximum interfacial area of 558 cm2. The compression rate was 3.75 cm2/min. External vibrations of the Langmuir trough were damped by an active vibration damper (Accurion).
Brewster Angle Microscopy. BAM images were obtained using a Brewster angle microscope (model BAM-300, KSV NIMA) coupled to a temperature-controlled Langmuir trough. A polarized He-Ne laser (wavelength 632.8 nm) was used as the light source, with the beam directed at the air/water interface at an angle of (the Brewster angle for water). The reflected beam was collected with a 10 objective; images were taken with a USB camera and corrected with a KSV NIMA software package. Measurements were performed at 21∘C and 23∘C.
Synchrotron measurements. GID and XSW studies were performed at the bending magnet beamline LANGMUIR (Kurchatov Center for Synchrotron Radiation, Moscow, Russia), which is equipped with a customized Langmuir trough. The closed Langmuir trough was flushed with water vapor saturated helium to reduce the scattering of the incident X-rays and to reduce the evaporation of the liquid subphase, thus keeping the liquid surface position fixed. The photon energy of the incident beam for all GID measurements was 13 keV. GID measurements were performed at ( is the critical angle of total external reflection for water). The diffracted intensity in the direction was recorded using a linear position sensitive detector (Mythen) with a vertical stripe size of m. For XSW measurements, the fluorescence signal was recorded by an energy dispersive Vortex EX detector mounted perpendicular to the liquid surface. The characteristic fluorescence spectra were recorded for each angle of incidence in the angular range corresponding to the total external reflection region.
@@ Acknowledgments
The authors are deeply indebted to Dr. Vladimir Kaganer, Dr. Boris Ostrovskii and Dr. Oleg Konovalov for fruitful discussions, insightful advice, and helpfulness.